# Images of black holes viewed by distant observer

### 03 Jan 2024 v1

@@ Appendix B Equations of motion for test particles
The first order differential equations of motion in the Kerr-Newman metric, derived by Brandon Carter Carter68 , are
$\Sigma\frac{dr}{d\tau}$ $=$ ±$\pm\sqrt{R(r)},$ (14)
$\Sigma\frac{d\theta}{d\tau}$ $=$ ±$\pm\sqrt{\Theta(\theta)},$ (15)
$\Sigma\frac{d\phi}{d\tau}$ $=$ $L\sin^{-2}\theta+a(\Delta^{-1}P-E),$ (16)
$\Sigma\frac{dt}{d\tau}$ $=$ $a(L-aE\sin^{2}\theta)+(r^{2}+a^{2})\Delta^{-1}P,$ (17)
@@ Appendix C Integral equations for test particle motion
For numerical calculations are very useful the integral equations of motion (14)–(17):
$\quad\>\>\>\fint\frac{dr}{\sqrt{R(r)}}=\fint\frac{d\theta}{\sqrt{\Theta(\theta)}},$ (18)
@@
The integrals in equations (18)–(21) monotonically grow along the particle trajectory and change sign at both the radial and polar turning points:
$\int^{r_{s}}_{r_{0}}\frac{dr}{\sqrt{R(r)}}=\int_{\theta_{0}}^{\theta_{s}}\frac{d\theta}{\sqrt{\Theta(\theta)}},$ (22)
where $r_{s}$ and $\theta_{s}$ are, respectively the initial and polar particle coordinates, and $r_{0}\gg r_{\rm h}$ and $\theta_{0}$ are the corresponding final particle coordinates. The more complicated case is with the one turning point in the latitude (or polar) direction, $\theta_{\rm min}(\lambda,q)$, which is a solution of the equation $\Theta(\theta)=0$. The corresponding ordinary integrals in equation (18) are written as
$\int_{r_{s}}^{r_{0}}\frac{dr}{\sqrt{R(r)}}=\int_{\theta_{\rm min}}^{\theta_{s}}\frac{d\theta}{\sqrt{\Theta(\theta)}}+\int_{\theta_{\rm min}}^{\theta_{0}}\frac{d\theta}{\sqrt{\Theta(\theta)}}.$ (23)

### 18 Jan 2024 v2

@@ Appendix B Equations of motion for test particles
The first order differential equations of motion in the Kerr-Newman metric, derived by Brandon Carter [4], are
$\Sigma\frac{dr}{d\tau}$ $=$ ±$\pm\sqrt{R(r)},$ (B.1)
$\Sigma\frac{d\theta}{d\tau}$ $=$ ±$\pm\sqrt{\Theta(\theta)},$ (B.2)
$\Sigma\frac{d\phi}{d\tau}$ $=$ $L\sin^{-2}\theta+a(\Delta^{-1}P-E),$ (B.3)
$\Sigma\frac{dt}{d\tau}$ $=$ $a(L-aE\sin^{2}\theta)+(r^{2}+a^{2})\Delta^{-1}P,$ (B.4)
@@ Appendix C Integral equations for test particle motion
For numerical calculations are very useful the integral equations of motion (B.1)–(B.4):
$\quad\>\>\>\fint\frac{dr}{\sqrt{R(r)}}=\fint\frac{d\theta}{\sqrt{\Theta(\theta)}},$ (C.1)
@@
The integrals in equations (C.1)–(C.4) monotonically grow along the particle trajectory and change sign at both the radial and polar turning points:
$\int^{r_{s}}_{r_{0}}\frac{dr}{\sqrt{R(r)}}=\int_{\theta_{0}}^{\theta_{s}}\frac{d\theta}{\sqrt{\Theta(\theta)}},$ (C.5)
where $r_{s}$ and $\theta_{s}$ are, respectively the initial and polar particle coordinates, and $r_{0}\gg r_{\rm h}$ and $\theta_{0}$ are the corresponding final particle coordinates. The more complicated case is with the one turning point in the latitude (or polar) direction, $\theta_{\rm min}(\lambda,q)$, which is a solution of the equation $\Theta(\theta)=0$. The corresponding ordinary integrals in equation (C.1) are written as
$\int_{r_{s}}^{r_{0}}\frac{dr}{\sqrt{R(r)}}=\int_{\theta_{\rm min}}^{\theta_{s}}\frac{d\theta}{\sqrt{\Theta(\theta)}}+\int_{\theta_{\rm min}}^{\theta_{0}}\frac{d\theta}{\sqrt{\Theta(\theta)}}.$ (C.6)